# Autonomous Driving in Unstructured Environments: How Far Have We Come?

### v3

@@ Abstract
Index Terms:
Autonomous Driving, Unstructured Environments, Comprehensive Survey.
In recent years, autonomous driving technology has wittnessed remarkable strides, particularly in structured environments such as urban areas and highways. Numerous autonomous driving systems have proven their capabilities in practical applications, thanks to their reliance on standardized markers like lane lines, traffic signals, and road signs. These markers enable autonomous vehicles to execute path planning and decision-making with substantial accuracy. However, the real world often presents environments that are not fully structured. As illustrated in Fig. 1, unstructured outdoor environments differ significantly from their structured urban counterparts. These unstructured environments like mountains, jungles, and deserts, lack clear road signs, traffic signals, and lane markings, and can include rural roads, off-road terrain, construction sites, parking lots, and other areas characterized by high complexity and unpredictability.
Figure 1: Differences between structured environments and unstructured environments are significant. Structured environments offer numerous structural cues, such as lane markings and roadway signs. In contrast, unstructured environments lack well-paved and clearly defined roadways.
@@
{forest}
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, where level=6font=, [ Autonomous Driving (AD) in Unstructured Environments [Modular AD, text width=5 em [Offline Mapping,text width=7.5 em [LiDAR-based, text width=7 em [Kaess et al. [1], Pan et al. [2], Song et al. [3], Lee et al. [4]…,leaf,text width=20.3em] ] [Fusion-based, text width=7 em [Ren et al. [5], Ren et al. [6], Dellaert et al. [7], Gao et al. [8]…,leaf,text width=20.3em] ] ] [Pose Estimation, text width=7.5 em [Matching-based, text width=7 em [Peng et al. [9], Li et al. [10], Ren et al. [6], Egger et al. [11]…,leaf,text width=20.3em] ] [Odometry-based, text width=7 em [Marks et al. [12], Zhang et al. [13], Yuan et al. [14], Chen et al. [15]…,leaf,text width=20.3em] ] ] [Environmental Perception,text width=7.5 em [Traversability Estimation, text width=7 em [LiDAR-based,text width=7 em [Thrun et al. [16], Chen et al. [17]…,leaf,text width=11.5 em] ] [Vision-based,text width=7 em [Gao et al. [18], Jung et al. [19]…,leaf,text width=11.5 em] ] [Fusion-based,text width=7 em [Yan et al. [20], Ye et al. [21]…,leaf,text width=11.5 em] ] ] [Semantic Segmentation, text width=7 em [LiDAR-based,text width=7 em [Liu et al. [22], Zhou et al. [23]…,leaf,text width=11.5 em] ] [Vision-based,text width=7 em [Singh et al. [24], Jiang et al. [25]…,leaf,text width=11.5 em] ] [Fusion-based,text width=7 em [Feng et al. [26], Kim and Seo [27]…,leaf,text width=11.5 em] ] ] ] [Path Planning,text width=7.5em [Global Path Planning,text width=7 em [Search-based, text width=7em [Stentz [28], Xu et al. [29]…,leaf,text width=11.5 em] ] [Sampling-based, text width=7em [Jaillet et al. [30], Jiang et al. [31]…,leaf,text width=11.5 em] ] [Biologically-inspired, text width=7em [Wang et al. [32], Zhu et al. [33]…,leaf,text width=11.5 em] ] ] [Local Path Planning,text width=7 em [Optimization-based, text width=7em [Lee and Wah [34], Guo et al. [35]…,leaf,text width=11.5 em] ] [Artificial Potential Field, text width=7em [Yao et al. [36], Jiang et al. [37]…,leaf,text width=11.5 em] ] [Dynamic Window, text width=7em [Yang et al. [38], Eirale et al. [39]…,leaf,text width=11.5 em] ] [Data-driven, text width=7em [Liu et al. [40], Wang et al. [41]…,leaf,text width=11.5 em] ] ] ] [Motion Control,text width=7.5em [PID Control, Preview Tracking Control, Sliding Mode Control, and MPC [42, 43]…, leaf,text width=29 em] ] ] [End-to-End AD, text width=5em [Imitation Learning,text width=7.5 em [Behavior Cloning: Leininger et al. [44] and Manderson et al. [45], Inverse RL: Zhu et al. [46]…, leaf,text width=29 em] ] [Reinforcement Learning,text width=7.5 em [Value-based, Policy-based: Liang et al. [47], Actor-critic: Weerakoon et al. [48]…, leaf,text width=29 em] ] ] [AD Dataset, text width=5 em [Traversability Estimation,text width=7.5 em [TrailNet [49], Vale [50], CaT [51], ORFD [52], Verti-Wheelers [53], M3-GMN [54]…, leaf,text width=29 em] ] [Semantic Segmentation,text width=7.5 em [DeepScene [55], YCOR [56], RUGD [57], RELLIS-3D [25], GOOSE [58], WildScenes [59]…, leaf,text width=29 em] ] [Other Datasets,text width=7.5 em [OPEDD [60], TartanDrive [61], Rally Estonia [62], Wild-Places [63]…, leaf,text width=29 em] ] ] ] ]
@@
Despite the challenges, autonomous driving in unstructured environments has attracted early research interest. For example, the 2004 DARPA Grand Challenge took place in a desert setting [64]. Some reviews [65] have summarized research in this area. Guastella and Muscato [66] focus on perception in autonomous driving within unstructured environments, while Borges et al. [67] and Islam et al. [68] examine traversability estimation in such contexts. Wijayathunga et al. [69] address both perception and path planning for autonomous driving in unstructured environments. The survey in [70] is dedicated to path planning, and Mortimer and Maehlisch [71] emphasize datasets relevant to autonomous driving in unstructured environments. However, existing reviews on autonomous driving in unstructured environments often focus on specific aspects such as traversability analysis, path planning, and datasets, while neglecting the holistic view of autonomous driving as an integrated system. Moreover, current reviews lack investigation into offline mapping and pose estimation for autonomous driving in unstructured environments. In this survey, we present a comprehensive review of autonomous driving in unstructured environments, as shown in Fig. 3.
@@
This review aims to provide a comprehensive overview and analysis of the current state, challenges, and future research directions in autonomous driving technology for unstructured environments. It is designed to offer valuable insights for newcomers entering this field and to foster critical discussion among established researchers. In Section I we pinpoint offline mapping in unstructured environments. In Section II we review pose estimation methods, followed by environmental perception methods in Section III. Section IV is dedicated to path planning methods, while motion control is discussed in Section V. End-to-end autonomous driving methods are discussed in Section VI. Section VII reviews available datasets related to autonomous driving in unstructured environments. In Section VIII we present several future research directions. Finally, in Section IX, we provide relevant conclusions about autonomous driving in unstructured environments.
@@ I Offline Mapping
I-A Challenges of Offline Mapping in Unstructured Environments
@@
The map construction process typically revolves around Simultaneous Localization and Mapping (SLAM) as the core technology, complemented by manual labeling and precise surveying techniques for both online and offline applications. Based on the types of sensors employed, the mapping methods for autonomous driving can be categorized into four main types: LiDAR-based methods, vision-based methods, millimeter-wave-based methods, and multi-sensor fusion methods, as shown in Table I. Thanks to their exceptional ranging accuracy and robust performance under varying lighting conditions, LiDAR-based mapping methods have seen extensive application in the field of map construction.
I-B LiDAR-based Methods
Wolcott and Eustice [82] construct a pose graph incorporating odometry, laser scan matching, and GPS prior constraints, solving the nonlinear least squares problem using the iSAM [1] method to optimize robot poses. Point clouds are reprojected into a global frame and accumulated into a global point cloud, stored via sparse histograms. The expectation-maximization algorithm is then applied to each histogram column to generate Gaussian mixture models for each grid cell, capturing z-height and reflectivity distributions, thus creating offline maps for online localization. Fu and Yu [73] propose a two-stage fusion algorithm for LiDAR scan matching in off-road environments, combining correlative scan matching (CSM) for large offsets and local scan matching techniques for smaller offsets, significantly enhancing accuracy and robustness in complex terrains.
Towards the large-scale unstructured environments, a robust off-line mapping system [5, 6] integrating GNSS and LiDAR data using a factor graph framework [7], introducing a new degeneration indicator for point cloud registration to maintain precision despite GNSS outages. Li et al. [10] utilize LiDAR and Global Navigation Satellite System (GNSS) during the offline mapping process to record tree trunk positions and collect point cloud data of the forest environment. The data is then processed to extract trunk points, project them onto a 2D plane, and perform Delaunay triangulation to create a global Delaunay Triangulation graph. This ensures an accurate global forest map for real-time localization of autonomous robots. For low-speed vehicles in unstructured environments, a mapping system [8] combining multiple sensors into a factor graph achieves globally consistent maps through robust two-stage optimization, correcting RTK measurements and detecting LiDAR odometry degeneration.
TABLE I: Offline mapping methods for unstructured environments.
@@ I-B LiDAR-based Methods
Wolcott and Eustice [82] LiDAR Mapping Multi-scene GPS/LiDAR
Fu and Yu [73] Point Clouds Registration Off-road LiDAR
Pan et al. [2] Dense Elevation Mapping KITTI/Campus LiDAR/Camera
Chahine et al. [83] Mapping Snowy Forest GPS/IMU/LiDAR/Camera
Ren et al. [5, 6] LiDAR Mapping Off-road GNSS/LiDAR
Gao et al. [8] LiDAR Mapping Residential Area RTK/IMU/LiDAR
Cui et al. [84] LiDAR Mapping Featureless Area RTK/IMU/LiDAR
Song et al. [3] Semantic Mapping Off-road GPS/IMU/LiDAR
Lee et al. [4] Online Mapping Unpaved Road GNSS/LiDAR/Camera
Zou et al. [85] LiDAR Mapping Multi-scene IMU/LiDAR
Lim [86] LiDAR Mapping Multi-scene LiDAR
Hu et al. [87] LiDAR Mapping Multi-scene LiDAR
@@
Long-term association LiDAR-IMU odometry systems, like LTA-OM [85], use corrected historical maps to provide global constraints during mapping, reducing drift and improving consistency over multiple sessions. Lim [86] presents the outlier-robust mapping techniques that leverage fast ground segmentation and graduated non-convexity (GNC) for handling significant data outliers, ensuring reliable mapping in dynamic environments. Finally, MS-Mapping [87] addresses data redundancy and pose graph scalability using Gaussian mixture model and Wasserstein distance-based keyframe selection, achieving high efficiency and accuracy in large-scale environments.
I-C Fusion-based Methods
@@ II Pose Estimation
II-A Challenges of Pose Estimation in Unstructured Environments
@@
II-B Matching-based Methods
TABLE II: Pose estimation methods using prior HD-Map for unstructured environments.
@@ II-B Matching-based Methods
Wolcott and Eustice [82] LiDAR Point Cloud Map Multi-scene IMU/LiDAR
Li et al. [10] Forest LiDAR
Ren et al. [6] Off-road IMU/LiDAR
Egger et al. [11] Multi-scene IMU/LiDAR
Zou et al. [85] Multi-scene IMU/LiDAR
Peng et al. [9] Changing Scene IMU/LiDAR
Hroob et al. [88] Changing Scene IMU/LiDAR
II-B1 Matching with High-Definition Map
For autonomous driving in unstructured environments, map-matching-based pose estimation method with High-Definition Map (HD-Map) is crucial for precise and safe navigation, as shown in Table II Some methods focus on the robustness and efficiency of localization within prior maps. Wolcott and Eustice [82] introduce a method utilizing multiresolution Gaussian mixture models (GMMs) to create maps capturing both three-dimensional structure and ground-plane reflectivity. This method allows for fast and accurate multiresolution inference, making real-time optimal localization feasible even under adverse conditions like heavy snowfall. The use of GPU processing further accelerates computation, enabling the system to handle all measured 3D points without spatial downsampling, thereby maintaining high localization accuracy.
Similarly, Li et al. [10] propose using Delaunay triangulation to match local point clouds to a global tree map for autonomous robots in forest environments. This approach achieves fast and real-time global localization, leveraging the Delaunay triangulation to solve vehicle tracking problems solely with LiDAR data, proving effective in complex forest environments. Additionally, Ren et al. [6] present a robust 3D offline mapping and online localization approach for off-road scenarios. The method includes a novel degeneracy indicator integrated into a factor graph framework, which helps detect and handle the degeneration state of scan matching algorithms. The system can seamlessly switch between map matching and LiDAR odometry modes, providing high accuracy and flexible localization solutions in challenging off-road conditions.
@@ II-B1 Matching with High-Definition Map
ROLL [9] incorporates temporary mapping to address long-term scene changes, and this system activates mapping processes when matching quality is low, discarding unreliable global matching poses and merging temporary key-frames into the global map once reliable association is achieved. Moreover, Hroob et al. [88] introduce a stability scan filter that learns and infers the motion status of objects, identifying stable landmarks for robust localization. Using an unsupervised labeling of LiDAR frames and a regression network, it effectively identifies long-term stable objects in agricultural environments, significantly improving localization performance by focusing on stable landmarks despite continuous environmental changes.
II-B2 Matching with Lightweight Map
TABLE III: Pose estimation methods using lightweight maps for unstructured environments.
@@ II-B2 Matching with Lightweight Map
Burger et al. [89] Road Network LiDAR
Ort et al. [90] Topometric Map LiDAR
Omama et al. [91] Topometric Map LiDAR/Camera
Si et al. [92] Topometric Map Camera
Ankenbauer et al. [93] Semantic Object Map LiDAR/Camera
In unstructured environments, while obtaining and maintaining HD-Map is costly, certain lightweight maps, such as topometric maps, are more cost-effective. These maps serve as efficient representations of the environment and can provide reliable and stable global constraints for the navigation and localization of autonomous vehicles [92], as shown in Table III.
One innovative method is presented by [89], which combines filter-based road course tracking and SLAM in unstructured rural areas lacking road markings or HD-Map. This approach employs a novel B-Spline measurement model tracked with an unscented Kalman filter, enhancing vehicle localization by correcting global position offsets through iterative closest point algorithms. Ort et al. [90] introduce MapLite, a system designed for autonomous intersection navigation without detailed prior maps or GPS localization. MapLite uses a sparse topometric map from OpenStreetMap [94] and integrates onboard sensors to segment the road, register the topometric map to sensor data, and plan trajectories using a variational path planner. This method successfully navigates over 15 km of rural roads with more than 100 intersection traversals, showcasing its effectiveness in unstructured environments.
Based on [90], Omama et al. [91] introduce ALT-Pilot, a system that utilizes publicly available road network information and crowdsourced language landmarks for navigation. ALT-Pilot employs a probabilistic multimodal localization algorithm integrating LiDAR, image, and language data, achieving high localization accuracy and robustness in both simulations and real-world scenarios. Similarly, Si et al. [92] propose TOM-odometry, a generalized localization framework that integrates odometry and topometric maps for GPS-denied scenarios. This framework employs odometry as a kinematic model for relative positioning, while the topometric map serves as a virtual global sensor to correct drift errors.
@@
II-B3 Place Recognition / Relocalization
TABLE IV: Place recognition methods for unstructured environments.

### v4

@@ Abstract
In recent years, autonomous driving technology has wittnessed remarkable strides, particularly in structured environments such as urban areas and highways. Numerous autonomous driving systems have proven their capabilities in practical applications, thanks to their reliance on standardized markers like lane lines, traffic signals, and road signs. These markers enable autonomous vehicles to execute path planning and decision-making with substantial accuracy. However, the real world often presents environments that are not fully structured. Unstructured outdoor environments differ significantly from their structured urban counterparts. These unstructured environments like mountains, jungles, and deserts, lack clear road signs, traffic signals, and lane markings, and can include rural roads, off-road terrain, construction sites, parking lots, and other areas characterized by high complexity and unpredictability.
@@
[…]
Autonomous Driving (AD) in Unstructured Environments
Modular AD Offline Mapping LiDAR-based [81], [148], … Fusion-based [157], [156], … Pose Estimation Matching-based [149], [105], … Odometry-based [121], [244], … Environmental Perception Traversability LiDAR-based [21] Vision-based [46] Fusion-based [234] Segmentation LiDAR-based [114] Vision-based [177] Fusion-based [38] Path Planning Global Search-based [179] Sampling-based [75] Biologically-inspired [199] Local Optimization-based [60] Artificial Potential Field [236] Dynamic Window [235] Data-driven [118] Motion Control PID, Preview Tracking Control, Sliding Mode Control, MPC, MPPI End-to-End AD Imitation Learning Behavior Cloning, Inverse Reinforcement Learning [256] Reinforcement Learning Value-based, Policy-based, Actor-critic [209], … AD Dataset Traversability Estimation ORFD [125], M3-GMN [223], … Semantic Segmentation DeepScene [191], YCOR [122], … Other Datasets OPEDD [137], TartanDrive [188], …
@@
Despite the challenges, autonomous driving in unstructured environments has attracted early research interest. For example, the 2004 DARPA Grand Challenge took place in a desert setting [33]. Some reviews [184] have summarized research in this area. [57] focus on perception in autonomous driving within unstructured environments, while [11] and [74] examine traversability estimation in such contexts. [215] address both perception and path planning for autonomous driving in unstructured environments. The survey in [203] is dedicated to path planning, and [134] emphasize datasets relevant to autonomous driving in unstructured environments. However, existing reviews on autonomous driving in unstructured environments often focus on specific aspects such as traversability analysis, path planning, and datasets, while neglecting the holistic view of autonomous driving as an integrated system. Moreover, current reviews lack investigation into offline mapping and pose estimation for autonomous driving in unstructured environments. In this survey, we present a comprehensive review of autonomous driving in unstructured environments, as shown in Fig. 2.
@@
This review aims to provide a comprehensive overview and analysis of the current state, challenges, and future research directions in autonomous driving technology for unstructured environments. It is designed to offer valuable insights for newcomers entering this field and to foster critical discussion among established researchers. In Section 1 we pinpoint offline mapping in unstructured environments. In Section 2 we review pose estimation methods, followed by environmental perception methods in Section 3. Section 4 is dedicated to path planning methods, while motion control is discussed in Section 5. End-to-end autonomous driving methods are discussed in Section 6. Section 8 reviews available datasets related to autonomous driving in unstructured environments. In Section 9 we present several future research directions. Finally, in Section 10, we provide relevant conclusions about autonomous driving in unstructured environments.
@@ 1 Offline Mapping
1.1 Challenges of Offline Mapping in Unstructured Environments
@@
The map construction process typically revolves around Simultaneous Localization and Mapping (SLAM) as the core technology, complemented by manual labeling and precise surveying techniques for both online and offline applications. Based on the types of sensors employed, the mapping methods for autonomous driving can be categorized into two main types: LiDAR-based methods and multi-sensor fusion methods, as shown in Table 1. Thanks to their exceptional ranging accuracy and robust performance under varying lighting conditions, LiDAR-based mapping methods have seen extensive application in the field of map construction.
1.2 LiDAR-based Methods
[217] construct a pose graph incorporating odometry, laser scan matching, and GPS prior constraints, solving the nonlinear least squares problem using the iSAM [81] method to optimize robot poses. Point clouds are reprojected into a global frame and accumulated into a global point cloud, stored via sparse histograms. The expectation-maximization algorithm is then applied to each histogram column to generate Gaussian mixture models for each grid cell, capturing z-height and reflectivity distributions, thus creating offline maps for online localization. [45] propose a two-stage fusion algorithm for LiDAR scan matching in off-road environments, combining correlative scan matching (CSM) for large offsets and local scan matching techniques for smaller offsets, significantly enhancing accuracy and robustness in complex terrains.
Towards the large-scale unstructured environments, a robust off-line mapping system [157, 156] integrating GNSS and LiDAR data using a factor graph framework [28], introducing a new degeneration indicator for point cloud registration to maintain precision despite GNSS outages. [105] utilize LiDAR and Global Navigation Satellite System (GNSS) during the offline mapping process to record tree trunk positions and collect point cloud data of the forest environment. The data is then processed to extract trunk points, project them onto a 2D plane, and perform Delaunay triangulation to create a global Delaunay Triangulation graph. This ensures an accurate global forest map for real-time localization of autonomous robots. For low-speed vehicles in unstructured environments, a mapping system [48] combining multiple sensors into a factor graph achieves globally consistent maps through robust two-stage optimization, correcting RTK measurements and detecting LiDAR odometry degeneration.
Table 1: Offline mapping methods for unstructured environments.
@@ 1.2 LiDAR-based Methods
[217] LiDAR Mapping Multi-scene GPS/LiDAR
[45] Point Clouds Registration Off-road LiDAR
[148] Dense Elevation Mapping KITTI/Campus LiDAR/Camera
[16] Mapping Snowy Forest GPS/IMU/LiDAR/Camera
[157, 156] LiDAR Mapping Off-road GNSS/LiDAR
[48] LiDAR Mapping Residential Area RTK/IMU/LiDAR
[24] LiDAR Mapping Featureless Area RTK/IMU/LiDAR
[178] Semantic Mapping Off-road GPS/IMU/LiDAR
[98] Online Mapping Unpaved Road GNSS/LiDAR/Camera
[259] LiDAR Mapping Multi-scene IMU/LiDAR
[111] LiDAR Mapping Multi-scene LiDAR
[70] LiDAR Mapping Multi-scene LiDAR
@@
Long-term association LiDAR-IMU odometry systems, like LTA-OM [259], use corrected historical maps to provide global constraints during mapping, reducing drift and improving consistency over multiple sessions. [111] presents the outlier-robust mapping techniques that leverage fast ground segmentation and graduated non-convexity (GNC) for handling significant data outliers, ensuring reliable mapping in dynamic environments. Finally, MS-Mapping [70] addresses data redundancy and pose graph scalability using Gaussian mixture model and Wasserstein distance-based keyframe selection, achieving high efficiency and accuracy in large-scale environments.
1.3 Fusion-based Methods
@@ 2 Pose Estimation
2.1 Challenges of Pose Estimation in Unstructured Environments
@@
2.2 Matching-based Methods
Table 2: Pose estimation methods using prior HD-Map for unstructured environments.
@@ 2.2 Matching-based Methods
[217] LiDAR Point Cloud Map Multi-scene IMU/LiDAR
[105] Forest LiDAR
[156] Off-road IMU/LiDAR
[34] Multi-scene IMU/LiDAR
[259] Multi-scene IMU/LiDAR
[149] Changing Scene IMU/LiDAR
[69] Changing Scene IMU/LiDAR
2.2.1 Matching with High-Definition Map
For autonomous driving in unstructured environments, map-matching-based pose estimation method with High-Definition Map (HD-Map) is crucial for precise and safe navigation, as shown in Table 2 Some methods focus on the robustness and efficiency of localization within prior maps. [217] introduce a method utilizing multiresolution Gaussian mixture models (GMMs) to create maps capturing both three-dimensional structure and ground-plane reflectivity. This method allows for fast and accurate multiresolution inference, making real-time optimal localization feasible even under adverse conditions like heavy snowfall. The use of GPU processing further accelerates computation, enabling the system to handle all measured 3D points without spatial downsampling, thereby maintaining high localization accuracy.
Similarly, [105] propose using Delaunay triangulation to match local point clouds to a global tree map for autonomous robots in forest environments. This approach achieves fast and real-time global localization, leveraging the Delaunay triangulation to solve vehicle tracking problems solely with LiDAR data, proving effective in complex forest environments. Additionally, [156] present a robust 3D offline mapping and online localization approach for off-road scenarios. The method includes a novel degeneracy indicator integrated into a factor graph framework, which helps detect and handle the degeneration state of scan matching algorithms. The system can seamlessly switch between map matching and LiDAR odometry modes, providing high accuracy and flexible localization solutions in challenging off-road conditions.
@@ 2.2.1 Matching with High-Definition Map
ROLL [149] incorporates temporary mapping to address long-term scene changes, and this system activates mapping processes when matching quality is low, discarding unreliable global matching poses and merging temporary key-frames into the global map once reliable association is achieved. Moreover, [69] introduce a stability scan filter that learns and infers the motion status of objects, identifying stable landmarks for robust localization. Using an unsupervised labeling of LiDAR frames and a regression network, it effectively identifies long-term stable objects in agricultural environments, significantly improving localization performance by focusing on stable landmarks despite continuous environmental changes.
2.2.2 Matching with Lightweight Map
Table 3: Pose estimation methods using lightweight maps for unstructured environments.
@@ 2.2.2 Matching with Lightweight Map
[12] Road Network LiDAR
[144] Topometric Map LiDAR
[142] Topometric Map LiDAR/Camera
[176] Topometric Map Camera
[4] Semantic Object Map LiDAR/Camera
In unstructured environments, while obtaining and maintaining HD-Map is costly, certain lightweight maps, such as topometric maps, are more cost-effective. These maps serve as efficient representations of the environment and can provide reliable and stable global constraints for the navigation and localization of autonomous vehicles [176], as shown in Table 3.
One innovative method is presented by [12], which combines filter-based road course tracking and SLAM in unstructured rural areas lacking road markings or HD-Map. This approach employs a novel B-Spline measurement model tracked with an unscented Kalman filter, enhancing vehicle localization by correcting global position offsets through iterative closest point algorithms. [144] introduce MapLite, a system designed for autonomous intersection navigation without detailed prior maps or GPS localization. MapLite uses a sparse topometric map from OpenStreetMap [62] and integrates onboard sensors to segment the road, register the topometric map to sensor data, and plan trajectories using a variational path planner. This method successfully navigates over 15 km of rural roads with more than 100 intersection traversals, showcasing its effectiveness in unstructured environments.
Based on [144], [142] introduce ALT-Pilot, a system that utilizes publicly available road network information and crowdsourced language landmarks for navigation. ALT-Pilot employs a probabilistic multimodal localization algorithm integrating LiDAR, image, and language data, achieving high localization accuracy and robustness in both simulations and real-world scenarios. Similarly, [176] propose TOM-odometry, a generalized localization framework that integrates odometry and topometric maps for GPS-denied scenarios. This framework employs odometry as a kinematic model for relative positioning, while the topometric map serves as a virtual global sensor to correct drift errors.
@@
2.2.3 Place Recognition / Relocalization
Table 4: Place recognition methods for unstructured environments.